For the confidence interval given by the formula X¯±1.19⋅σn−−√, what is the corresponding confidence level? For the purposes of this question, we can assume that n is sufficiently large, and σ is known.

Express your answer as a percent, but do NOT include the percent sign (%) in your response.

Round your response to at least 2 decimal places.
Answer #1

Solution :

Z\alpha/2 = 1.19

\alpha/2 = 0.117

\alpha = 0.117 * 2 = 0.234

c = 1 - \alpha = 1 - 0.234 = 0.766

Confidence level = 76.60